# Review of Demographic Bias in Face Recognition

### v1

@@ V-B In-Processing Methods
Amini et al. [87] proposed a debiasing algorithm that adjusts the sampling probabilities of data points in large datasets to reduce hidden biases. When applied for face detection use-case, their algorithm led to decrease in race and gender bias while improving classification accuracy. To our knowledge, similar approaches have not been tested for recognition or verification applications. In [53], Wang and Deng introduced a reinforcement learning-based race balance network (RL-RBN) where they applied adaptive margins through deep Q-learning. Their method aimed to reduce the skewness of feature scatter between racial groups, leading to more balanced performance across different demographics. As a part of this work, they also released two datasets- BUPT-GlobalFace and BUPT-BalancedFace datasets- that were specifically designed to study racial bias in FR systems.
Gong et al. proposed a group-adaptive classifier (GAC) that uses adaptive convolution kernels and attention mechanisms tailored to different demographic groups [88]. By applying kernel masks and attention maps specific to each group, their method activates facial regions that are more discriminative for each demographic, thereby improving recognition accuracy and fairness across demographic groups. Another approach by the same authors [74] introduced an adversarial network, called DebFace, which utilizes a multi-task learning framework to simultaneously learn identity and demographic attributes. Their method employed adversarial training to disentangle identity features from demographic attributes such as gender, age, and race to effectively reducing bias in the recognition process. Their experiments demonstrated that, for DebFace, not only recognition but also demographic attribute estimation tasks were less biased.
Reference Year Type of Method Test Dataset Method Description
@@
Gong et al. [74] 2020 In-Processing RFW Adversarial debiasing network with identity and demographic classifiers
@@
Terhörst et al. [96] 2020 Post-Processing Adience, Color-Feret, MORPH Fair score normalization to mitigate demographic bias
@@
TABLE IV: Summary of recent works on mitigation of demographic bias in face recognition.
@@
Another approach for domain-specific bias mitigation using disentangled representation learning was proposed by Liang et al. [89]. They introduced a two-stage method combining modules for disentangled representation learning with additive adversarial learning (AAL). While this work does not directly address demographic bias, it provides useful insights into how domain-specific biases can be mitigated by learning disentangled representations. The effectiveness of this method in reducing bias across various domains suggests its potential applicability in the context of demographic bias in FR. In [90], a Progressive Cross Transformer (PCT) was proposed to mitigate racial bias by decoupling face representations into identity-related and race-induced components. Using dual cross-transformers, the PCT refines identity features and suppresses racial noise, demonstrating lower racial bias without compromising recognition accuracy. The concept of score normalization was incorporated as a regularization term into the training objective enabling simultaneous optimization of recognition accuracy and demographic fairness [75]. This was facilitated by constraining the output scores of mated and non-mated pairs to adhere to a pre-defined distribution, and followed by minimizing differences in score distributions across demographic groups. During inference, the overall pipeline did not require modifications as the FR CNN architecture remained unaltered while only the weights were fine-tuned to the new objective.
@@
Finally, we discuss a couple of works dealing with fairness in facial attribute recognition. Both of these are based on contrastive learning– which can be useful mechanism to address the demographic bias in FR as well. Park et al. addressed fairness issues in attribute classification using a contrastive learning framework [91]. They constructed a Fair Supervised Contrastive Loss (FSCL) which reduces bias by normalizing intra-group compactness and inter-group separability, penalizing sensitive attribute information in representations. Although predominantly an in-processing method, the authors also incorporate a loss function to address the imbalance in training data where majority groups are constrained to have a better intra- group compactness and inter-class separability compared to the under-represented ones. Similarly, Zhang et al. proposed Fairness-aware Contrastive Learning (FairCL) for unsupervised representation learning and demonstrated the use-case of facial attribute recognition [92]. In addition to fair contrastive learning of feature representations, they also attempted to address the dataset bias by specifically generating contrastive sample pairs that share the same visual information apart from sensitive attributes. They also suggested unsupervised feature reweighting to strike balance the utility and fairness of learned representations.
@@ V-C Post-Processing Methods
These techniques are applied after the feature extraction and matching processes to adjust the final decision scores and reduce bias. While these methods are less commonly used compared to pre- and in-processing techniques, they can still play a role in ensuring equitable outcomes. Additionally, these approaches are relatively easy to integrate into existing systems.
@@
Michalski et al. investigated the impact of age variation on FR, particularly for children [93]. They showed that dynamic thresholding improves performance. To address age-related bias, they adjusted thresholds based on age differences as opposed to a fixed threshold. In another work on age-related bias, Srinivas et al. experimented with score-level fusion strategies to improve recognition accuracy for the children (age as demographic) [94]. They considered six fusion schemes that combined different score-normalization techniques and fusion rules. For normalizing, they considered z-norm and min-max strategies; while fusion was conducted using min, max, or sum rules. Robinson et al. [57] showed that applying a single threshold across different demographic groups leads to significant variations in the FMR. They addressed this issue by using per-subgroup thresholds to balance the FMRs across ethnic and gender groups, improving both recognition fairness and performance.
@@
A typical FR pipeline employs similarity functions to obtain a matching score. Terhorst et al. replaced the conventional similarity function by a fairness-driven neural network classifier [95]. By adding a penalization term in the loss function, their method was able to equalizes score distributions across ethnic groups, reducing intra-ethnic bias while maintaining high recognition performance. In another work [96], Terhorst et al. introduced an unsupervised fair score normalization method based on individual fairness principles, which treats similar individuals similarly. During training, they partitioned the identities in finite number of groups using K-means clustering on face embeddings. At inference, they computed the cluster-specific thresholds for both samples contributing to the score, and these threshold were combined with a global threshold to yield normalized scores. In a recent work, Linghu et al. extended traditional score normalization methods, such as Z and T normalization, by incorporating demographic information to enhance fairness in FR systems [97]. Furthermore, they evaluated three cohort-based approaches based on imposter scores, Platt scaling, and bi-modal cumulative distribution functions (CDF). Their findings demonstrated that the proposed method improved fairness across both race and gender demographic groups, particularly at low FMRs.
The works on mitigation are summarized in Table IV.
@@ VI Future Directions
In recent years, with advanced architectures and increase in the number of layers and parameters, FR models have gained a substantial improvement in their capacity to learn complex facial representations. These deeper architectures, often comprising over a hundred layers with millions of parameters, have significantly enhanced the ability of FR systems to generalize across challenging scenarios, resulting in higher overall recognition performance (as well as reduced bias). Additionally, the availability of larger, more diverse datasets has contributed to better learning outcomes. These datasets, which incorporate substantial variations demographics and covariates such as pose, illumination, and expression (PIE), have facilitated measurable progress in improving both accuracy and fairness FR.
@@
In this section, we examine some of the emerging challenges associated with demographic bias. These challenges highlight the need for ongoing research to adapt bias mitigation strategies to align with the advancements in FR applications.
@@
Bias in Lightweight Models: The lightweight FR models, often used in handheld devices and resource-constrained environments, encounter significant challenges concerning demographic bias. These systems, crucial for privacy-sensitive applications, often inherit limitations in capacity and architecture, leading to non-equitable performance across demographic groups. Bias in lightweight models has garnered limited research attention, despite their widespread deployment in mobile and IoT devices with varying sensor qualities. Techniques like knowledge distillation (KD) and pruning, while essential for model compression, introduce or amplify bias. For instance, Liu et al. highlighted that KD inherits biases from larger teacher models [98], while pruning has been shown to disproportionately impacts underrepresented groups [99, 100]. Incorporating fairness-aware techniques is crucial for mitigating these issues. Lin et al. [101] introduced FairGRAPE, a pruning method that evaluates network connections with demographic considerations, reducing performance disparities. Achieving demographic fairness in lightweight models requires targeted compression strategies and ethical evaluations of demographic-specific impacts. By integrating fairness principles into compression techniques, lightweight FR systems can achieve more equitable outcomes while maintaining efficiency and accuracy.
@@
Bias in Quantization: Quantization, a model compression technique, retains the original architecture (as opposed to KD and pruning– which often modify the structure) while reducing parameter precision, producing smaller and faster models. However, it can lead to higher demographic bias by prioritizing global performance over the accurate classification of under-represented groups. Such bias underscores the need for rigorous fairness evaluation across demographic subgroups when deploying compressed models. Quantization converts floating-point (FP) models into lower-precision formats like 8-bit, balancing efficiency and accuracy [102]. It typically consists of two approaches: post-training quantization (PTQ) and quantization-aware training (QAT). Some studies, such as Stoychev et al. demonstrated that 8-bit PTQ maintained fairness and accuracy in gender bias for face expression recognition [102]. However, reducing precision to 6 bits significantly degraded fairness, indicating a trade-off between compression and bias mitigation. Although similar investigations for FR systems remain limited, these findings highlight the need to evaluate and ensure demographic fairness in quantized models.
@@
Low Resolution: The research on demographic bias in FR has predominantly focused on high-resolution images, often overlooking the challenges posed by low-resolution images typically captured by surveillance cameras or from significant distances. One of the primary impediment to research this issue is the scarcity of low-resolution datasets that include demographic attributes. Consequently, demographic disparities in low-resolution FR remain under-explored, despite their importance in real-world applications. A recent work from Atzori et al. attempted to address this gap by designing a novel framework to investigate demographic bias in low-resolution FR [103]. They trained state-of-the-art FR models on various combinations of high- and low-resolution images. Testing on degraded images from five datasets revealed significant disparities across gender and ethnic groups, underscoring the need for timely interventions in low-resolution FR. It may be noted that their approach involved use of a generative model to convert high-resolution face images into realistic low-resolution counterparts. The importance of low-resolution FR is evident in programs like BRIAR [^5] BRIAR Programme by IARPA., which aim to enhance recognition technologies for challenging scenarios, such as long-distance identification and low-quality image acquisition. To tackle demographic bias in low-resolution FR, there is a need to develop both datasets and models tailored to these unique use-cases.
@@
Training Datasets: Demographically balanced datasets, while not entirely eliminating bias, significantly contribute to reducing non-equitable performance in FR. Thus, large, diverse, and balanced datasets are pivotal for achieving fairer and accurate models. However, acquiring such datasets is increasingly challenging due to cost and ethical and privacy concerns surrounding biometric data collection. The commercial sector, in particular, faces difficulties as most available datasets are collected locally from consenting individuals— which are often in limited size and demographic representation. These constraints necessitate innovative approaches to address demographic bias using smaller or synthetic datasets.
@@
The use of synthetic data in FR has recently gained traction as a potential solution to privacy and data-sharing concerns. Competitions like the FRSyn series have encouraged advancements in synthetic data usage [77, 78, 76]. Despite these efforts, FR models trained exclusively on synthetic datasets continue to underperform compared to those trained on real datasets of similar size [76, 104]. This gap is evident in both recognition accuracy (measured by metrics like FMR and FNMR) and demographic fairness (assessed by standard deviation of performance across groups). The analysis of synthetic datasets by Huber et al. revealed that demographic bias might worsen compared to the (real) training dataset [105]. Enhancing the quality and utility of synthetic datasets, beyond the aspect of bias, remains an open problem, requiring further exploration.
@@
Recognizing the increasing reliance on remote verification, the MdTF and DHS S&T introduced the Remote Identity Validation Technology Demonstration (RIVTD) initiative [^6] Remote Identity Validation Technology Demonstration (RIVTD). In addition to security, accuracy, and liveness detection requirements, this program also places particular emphasis on ensuring demographic fairness in such technologies. A recent study by Fatima et al. [106] investigated demographic fairness in RIdV technologies using statistical methods to analyze performance disparities. Their analysis of five commercial RIdV systems revealed that only two achieved equitable outcomes across demographic groups. Notably, higher FNMRs were observed among African cohorts and individuals with darker skin-tones. Such findings highlight the necessity of evaluating RIdV technologies across demographic groups to ensure equitable and unbiased performance.
@@
Complex Bias Factors (Intersectionality): The majority of research on mitigation of bias in FR, as discussed in Sec V has focused on single demographic attribute, such as race, age, or gender. However, several studies have identified that combination or intersection of various demographic factors causes (or amplifies) bias in FR models (cf. Sec II). Existing bias mitigation techniques typically target one demographic attribute at a time, achieving measurable improvements in fairness for that specific attribute. However, it remains unclear whether such processing inadvertently introduces imbalances in other demographic attributes. For instance, enhancing fairness for gender-related bias may increase disparities linked to ethnicity and vice-versa. This highlights the need for systematic evaluations of the intersectionality of demographic factors, such as race and gender combined. Consequently, developing mitigation methods capable of addressing multiple demographic attributes simultaneously remains an open challenge.
@@
Noisy Labels: The assignment of demographic attributes such as race, ethnicity, and skin-tone in FR datasets typically involves discrete labeling into finite categories. Some attributes, such as race, are often self-reported. In many cases, race and ethnicity annotations may be derived from automatic classifiers or manual efforts. Automatic classifiers, predominantly based on deep learning, are likely to be susceptible to bias too; while manual annotations are prone to human judgment errors. Similarly, skin-tone is frequently categorized using scales like Fitzpatrick’s, which discretizes it into specific values, ignoring its continuous spectrum. This reliance on discrete labels introduces noise into the training data, as samples near category boundaries are often inaccurately labeled. Recent work [86], addressed this issue by using probabilistic weights (soft labels) for demographic information instead of utilizing rigif (categorical) labels. However, most existing methods overlook the issue of errors in training data.

### v2

@@ V-B In-Processing Methods
Amini et al. [109] proposed a debiasing algorithm that adjusts the sampling probabilities of data points in large datasets to reduce hidden biases. In the face detection use-case, their algorithm led to a decrease in race and gender bias while improving classification accuracy. To our knowledge, similar approaches have not been tested for recognition or verification applications. In [73], Wang and Deng introduced a reinforcement learning-based race balance network (RL-RBN) where they applied adaptive margins through deep Q-learning. Their method aimed to reduce the skewness of feature scatter between racial groups, leading to more balanced performance across different demographics. As a part of this work, they also released two datasets— BUPT-GlobalFace and BUPT-BalancedFace datasets— that were specifically designed to study racial bias in FR systems.
Gong et al. proposed a group-adaptive classifier (GAC) that uses adaptive convolution kernels and attention mechanisms tailored to different demographic groups [110]. By applying kernel masks and attention maps specific to each group, their method activates facial regions that are more discriminative for each demographic, thereby improving recognition accuracy and fairness across demographic groups. Another approach by the same authors [95] introduced an adversarial network, called DebFace, which utilizes a multi-task learning framework to simultaneously learn identity and demographic attributes. Their method employed adversarial training to disentangle identity features from demographic attributes such as gender, age, and race to effectively reducing differential in the recognition process. Their experiments demonstrated that, for DebFace, not only recognition but also demographic attribute estimation tasks were less biased. In [111], authors proposed unsupervised data partitioning, in iterative manner, followed by Invariant Feature Regularization, to improve generalization across demographic groups. The unsupervised partitioning was chosen as an inexpensive alternative to annotating training data with multiple (confounding) attributes.
TABLE V: Summary of recent works on mitigation of demographic bias in face recognition.
Reference Year Type of Method Test Dataset Summary
@@
Yucer et al. [105] 2020 Data-Processing RFW Adversarial image-to-image augmentation
@@
Gong et al. [95] 2020 In-Processing RFW Adversarial debiasing with identity and demographic classifiers
@@
Ma et al. [111] 2023 In-Processing MFR, CelebA, RFW Unsupervised data partitioning and Invariant Feature Regularization
@@
Xu et al.  [116] 2021 In-Processing RFW, BFW Additional penalty for disparities in FPR across instances
Li et al.  [117] 2025 In-Processing RFW, BFW, NFW Customized instance margins to preserve consistency in FPR and TPR during the label classification in training
@@
Morales et al. [74] 2020 Post-Processing DiveFace, CelebA Disentangling sensitive attributes from features for privacy preservation
Terhörst et al. [121] 2020 Post-Processing Color-Feret, MORPH Fair score normalization to mitigate demographic bias
Dhar et al. [122] 2021 Post-Processing IJB-C Privacy-preserving disentanglement of features via adversarial training
Salvador et al. [123] 2021 Post-Processing RFW, BFW Conditional calibration of clustered embeddings
Conti et al. [124] 2022 Post-Processing LFW Employing shallow MLP on the FR embeddings
Robinson et al. [76] 2023 Post-Processing BFW Privacy preserving domain adaptation framework reproject embeddings
@@
Liu et al. [126] 2022 Post-Processing RFW Threshold Consistency Penalty (TCP) to enforce consistent decision boundaries across multiple datasets
@@
Another approach for domain-specific bias mitigation using disentangled representation learning was proposed by Liang et al. [112]. They introduced a two-stage method combining modules for disentangled representation learning with additive adversarial learning (AAL). While this work does not directly address accuracy differential, it provides useful insights into how domain-specific biases can be mitigated by learning disentangled representations. The effectiveness of this method in reducing bias across various domains suggests its potential applicability in the context of demographic fairness in FR. In [113], a Progressive Cross Transformer (PCT) was proposed to mitigate racial bias by decoupling face representations into identity-related and race-induced components. Using dual cross-transformers, the PCT refines identity features and suppresses racial noise, demonstrating lower racial differentials without compromising recognition accuracy. The concept of score normalization was incorporated as a regularization term into the training objective enabling simultaneous optimization of recognition accuracy and demographic fairness [96]. This was facilitated by constraining the output scores of mated and non-mated pairs to adhere to a pre-defined distribution, and followed by minimizing differences in score distributions across demographic groups. During inference, the overall pipeline did not require modifications as the FR CNN architecture remained unaltered while only the weights were fine-tuned to the new objective.
A couple of works [116, 117] have proposed customized instance margins aimed at preserving consistent error rates– either false positives or false positives combined with true positives. These methods modify the Softmax-based loss by additional penalty terms derived from disparities in the concerned error rates.
@@
Finally, we discuss a couple of works dealing with fairness in facial attribute recognition. Both of these are based on contrastive learning– which can be useful mechanism to address the demographic fairness in FR as well. Park et al. addressed fairness issues in attribute classification using a contrastive learning framework [114]. They constructed a Fair Supervised Contrastive Loss (FSCL) which reduces disparities by normalizing intra-group compactness and inter-group separability, penalizing sensitive attribute information in representations. Although predominantly an in-processing method, the authors also incorporate a loss function to address the imbalance in training data where majority groups are constrained to have a better intra- group compactness and inter-class separability compared to the under-represented ones. Similarly, Zhang et al. proposed Fairness-aware Contrastive Learning (FairCL) for unsupervised representation learning and demonstrated the use-case of facial attribute recognition [115]. In addition to fair contrastive learning of feature representations, they also attempted to address the dataset bias by specifically generating contrastive sample pairs that share the same visual information apart from sensitive attributes. They also suggested unsupervised feature reweighting to strike balance between the utility and fairness of learned representations.
@@ V-C Post-Processing Methods
These techniques are applied after the feature extraction and/or matching processes to adjust the final decision scores and reduce bias. We also consider methods that are devised to process features (embeddings) extracted from a pretrained or frozen FR network. While these methods are less commonly used compared to pre- and in-processing techniques, they can still play a role in ensuring equitable outcomes. Additionally, these approaches are relatively easy to integrate into existing systems.
@@
Michalski et al. investigated the impact of age variation on FR, particularly for children [118]. They showed that dynamic thresholding improves performance. To address age-related accuracy differences, they adjusted thresholds based on age differences as opposed to a fixed threshold. In another work on age-related bias, Srinivas et al. experimented with score-level fusion strategies to improve recognition accuracy for the children (age as demographic) [119]. They considered six fusion schemes that combined different score-normalization techniques and fusion rules. For normalization, they considered $\mathbb{Z}$-norm and $\min$-$\max$ strategies; while fusion was conducted using min, max, or sum rules. Robinson et al. [77] showed that applying a single threshold across different demographic groups leads to significant variations in the FMR. They addressed this issue by using per-subgroup thresholds to balance the FMRs across ethnic and gender groups, improving both recognition fairness and performance.
@@
A typical FR pipeline employs similarity functions to obtain a matching score. Terhörst et al. replaced the conventional similarity function by a fairness-driven neural network classifier [120]. By adding a penalization term in the loss function, their method was able to equalizes score distributions across ethnic groups, reducing intra-ethnic bias while maintaining high recognition performance. In another work [121], Terhörst et al. introduced an unsupervised fair score normalization method based on individual fairness principles, which treats similar individuals similarly. During training, they partitioned the identities in finite number of groups using K-means clustering on face embeddings. At inference, they computed the cluster-specific thresholds for both samples contributing to the score, and these threshold were combined with a global threshold to yield normalized scores. Morales et al. [74] introduced SensitiveNets, a framework that enforces the removal of sensitive attribute information from learned feature representations. By disentangling sensitive attributes from task-relevant features, their method ensures that downstream agents cannot exploit protected information, thereby enhancing both privacy and fairness.
Salvador et al. proposed FairCal: a post-processing method aimed at improving fairness in face verification without retraining or requiring sensitive attribute labels [123]. They clustered image embeddings (obtained from frozen FR network) in an unsupervised manner and applied conditional calibration maps per cluster. They reported reduced false matches while improving overall accuracy. In [124], Conti et al. employed a shallow MLP on top of the frozen FR model as a post-processing strategy to mitigate gender bias in FR. They enhanced the representations of underperforming subgroups using a Fair von Mises-Fisher loss with class-specific variance control. In a recent work, Linghu et al. extended traditional score normalization methods, such as $\mathbb{Z}$ and $\mathbb{T}$ normalization, by incorporating demographic information to enhance fairness in FR systems [125]. Furthermore, they evaluated three cohort-based approaches based on imposter scores, Platt scaling, and bi-modal cumulative distribution functions (CDF). Their findings demonstrated that the proposed method improved fairness across both race and gender demographic groups, particularly at low FMRs.
Liu et al. [126] argued that the existing FR evaluation protocols, which rely on dataset-specific thresholds, are practical, and they introduced the One-Threshold-for-All (OTA) protocol that employs a single calibration threshold across all datasets. To optimize performance under the OTA setting, they also proposed the Threshold Consistency Penalty (TCP), which enforces consistent decision boundaries across domains through implicit domain division followed by calibration and estimation of domain thresholds.
Feature descriptors, or embeddings, generated by FR networks inherently encode sensitive identity-related attributes, such as gender and skin-tone; thereby rendering them vulnerable to privacy risks. Although the preservation of privacy lies outside the scope of this review, it is closely linked to concerns of fairness, given its role in exacerbating performance disparities across demographic groups. Research on disentangling sensitive attributes from identity-related representations remains relatively sparse. In [127], Dhar et al. proposed the adversarial gender de-biasing algorithm, which aims to reduce gender-specific information from embeddings extracted by pre-trained (frozen) FR networks. This method employs adversarial training to simultaneously suppress gender information and maintain identity recognition performance. Building on this approach, the Protected Attribute Suppression System (PASS) [122] advances the concept by introducing a descriptor-level adversarial framework to mitigate accuracy differentials. PASS operates directly on pre-trained embeddings and employs a novel discriminator-based technique to suppress attributes such as gender and skin tone without necessitating end-to-end retraining. Additionally, Robinson et al. [76] highlighted that applying a single decision threshold across demographic subgroups results in disparate performance, diverging from global evaluation metrics. To address this, they proposed a domain adaptation framework that re-projects embeddings into a feature space where demographic attributes are suppressed, while identity-relevant information is retained.
The works on mitigation are summarized in Table V. For most of the state-of-the-art works aimed towards mitigating accuracy differentials, the FR models have been trained on massive datasets, fine-tuned primarily for accuracy, and are potentially operating near their capacity limits. When a new objective, such as fairness, is introduced, it competes with the existing accuracy objective, often resulting in performance degradation. Some recent methods have demonstrated the potential to improve fairness metrics without sacrificing overall accuracy. These methods, however, often involve architectural modifications, such as the addition of auxiliary modules. It is worth noting that such models, by virtue of their altered structure, differ from their original baselines (with more learnable parameters), and thus may not be directly comparable.
@@ VI Future Directions
In recent years, with advanced architectures and increase in the number of layers and parameters, FR models have gained a substantial improvement in their capacity to learn complex facial representations. These deeper architectures, often comprising over a hundred layers with millions of parameters, have significantly enhanced the ability of FR systems to generalize across challenging scenarios, resulting in higher overall recognition performance (as well as improved fairness). Additionally, the availability of larger, more diverse datasets has contributed to better learning outcomes. These datasets, which incorporate substantial variation in demographics and covariates such as pose, illumination, and expression (PIE), have facilitated measurable progress in improving both accuracy and fairness FR.
@@
In this section, we examine some of the emerging challenges associated with demographic fairness. These challenges highlight the need for ongoing research to adapt bias mitigation strategies to align with the advancements in FR applications.
@@
Lightweight Models: The lightweight FR models, often used in handheld devices and resource-constrained environments, encounter significant challenges concerning demographic fairness. These systems, crucial for privacy-sensitive applications, often inherit limitations in capacity and architecture, leading to non-equitable performance across demographic groups. Bias in lightweight models has garnered limited research attention, despite their widespread deployment in mobile and IoT devices with varying sensor qualities. Techniques like knowledge distillation (KD) and pruning, while essential for model compression, introduce or amplify bias. For instance, Liu et al. highlighted that KD inherits biases from larger teacher models [128], while pruning has been shown to disproportionately impacts underrepresented groups [129, 130]. Incorporating fairness-aware techniques is crucial for mitigating these issues. Lin et al. [131] introduced FairGRAPE, a pruning method that evaluates network connections with demographic considerations, reducing performance disparities. Caldeira et al. [132] proposed a multi-specialized teacher framework where each teacher model is trained on a specific ethnicity and collectively distills knowledge into a student model. Achieving demographic fairness in lightweight models requires targeted compression strategies and ethical evaluations of demographic-specific impacts. By integrating fairness principles into compression techniques, lightweight FR systems can achieve more equitable outcomes while maintaining efficiency and accuracy.
@@
Quantization of Models and Data: Quantization, a model compression technique, retains the original architecture (as opposed to KD and pruning– which often modify the structure) while reducing parameter precision, producing smaller and faster models. However, it can lead to higher demographic bias by prioritizing global performance over the accurate classification of under-represented groups. Such bias underscores the need for rigorous fairness evaluation across demographic subgroups when deploying compressed models. Quantization converts floating-point (FP) models into lower-precision formats like 8-bit, balancing efficiency and accuracy [133]. It typically consists of two approaches: post-training quantization (PTQ) and quantization-aware training (QAT). Some studies, such as Stoychev et al. demonstrated that 8-bit PTQ maintained fairness and accuracy in gender bias for face expression recognition [133]. However, reducing precision to 6 bits significantly degraded fairness, indicating a trade-off between compression and bias mitigation. Although similar investigations for FR systems remain limited, these findings highlight the need to evaluate and ensure demographic fairness in quantized models.
While the majority of research in this area focuses on quantized models, a number of recent studies have investigated the implications of lossy image compression on fairness in FR. The work in [134] explored how lossy JPEG compression influences recognition accuracy across demographic groups, more specifically on racial phenotypes. Their findings indicate that compression disproportionately degrades performance for certain traits, such as darker skin-tones and wider facial features. They also demonstrated that disabling chroma sub-sampling can improve FMRs across several racial phenotypes. Similarly, Qiu et al. [135] examined the effects of five neural image compression methods and found that certain phenotypic features–such as skin-tone and hair-type–are more susceptible to degradation at lower bit rates. Notably, at extremely low bit rates, the amplification of bias was particularly pronounced for individuals belonging to African race.
@@
Low Resolution: The research on demographic fairness in FR has predominantly focused on high-resolution images, often overlooking the challenges posed by low-resolution images typically captured by surveillance cameras or from significant distances. One of the primary impediment to research this issue is the scarcity of low-resolution datasets that include demographic attributes. Consequently, demographic disparities in low-resolution FR remain under-explored, despite their importance in real-world applications. A recent work from Atzori et al. attempted to address this gap by designing a novel framework to investigate demographic bias in low-resolution FR [136]. They trained state-of-the-art FR models on various combinations of high- and low-resolution images. Testing on degraded images from five datasets revealed significant disparities across gender and ethnic groups, underscoring the need for timely interventions in low-resolution FR. It may be noted that their approach involved use of a generative model to convert high-resolution face images into realistic low-resolution counterparts. In [45, 59], the impact of blur and image resolution were investigated in the context of 1:$N$ identification, revealing that increased blurring leads to elevated FPIRs. These studies also noted that the impact of blur varies across demographic attributes, particularly race and gender. The quality degradations of blur and resolution in these studies were simulated using Gaussian blurring and bicubic interpolation, respectively. The importance of low-resolution FR is evident in programs like BRIAR [^3] BRIAR Programme by IARPA., which aim to enhance recognition technologies for challenging scenarios, such as long-distance identification and low-quality image acquisition. To tackle demographic fairness in low-resolution FR, there is a need to develop both datasets and models tailored to these unique use-cases.
@@
Datasets for Training and Evaluation: While demographically balanced datasets alone may not fully eliminate bias, they play a critical role in reducing disparities in FR performance. Achieving fairness requires not only demographic diversity (e.g., across race, gender, and age) but also balance with respect to non-demographic or soft attributes such as hairstyle, facial hair, makeup, and occlusion. Therefore, large-scale datasets that are balanced across both demographic and non-demographic dimensions are essential for developing FR models that are both fair and accurate. However, acquiring such datasets is increasingly challenging due to cost and ethical and privacy concerns surrounding biometric data collection. The commercial sector, in particular, faces difficulties as most available datasets are collected locally from consenting individuals—which are often in limited size and demographic representation. These constraints necessitate innovative approaches to address demographic fairness using smaller or synthetic datasets.
Access to datasets annotated with multiple attributes—both demographic and non-demographic—will prove to be a valuable resource for advancing fairness in biometric recognition. Such datasets enable deeper exploration of the underlying causes of bias by facilitating the analysis of correlations and potential causal relationships among factors that influence recognition performance. Moreover, they provide a foundation for designing effective mitigation strategies. When used for assessment, multi-attribute datasets can help controlling or isolating individual factors, enabling systematic benchmarking of their specific impact on model performance.
@@
The use of synthetic data in FR has recently gained traction as a potential solution to privacy and data-sharing concerns. Competitions like the FRSyn series have encouraged advancements in synthetic data usage [98, 99, 97]. Despite these efforts, FR models trained exclusively on synthetic datasets continue to underperform compared to those trained on real datasets of similar size [97, 137]. This gap is evident in both recognition accuracy (measured by metrics like FMR and FNMR) and demographic fairness (assessed by standard deviation of performance across groups). The analysis of synthetic datasets by Huber et al. revealed that demographic bias might worsen compared to the (real) training dataset [138]. Enhancing the quality and utility of synthetic datasets, beyond the aspect of fairness, remains an open problem, requiring further exploration.
Causes of Bias: Recent research indicates that accuracy disparities across demographic groups in FR systems are not solely attributable to inherent physical or biological characteristics. Instead, these disparities are frequently influenced by non-demographic factors– such as hairstyle, cosmetic use, or occlusion patterns– that are shaped by cultural norms, social behaviors, or individual preferences. Despite their relevance, such factors have historically received limited attention, with only a few recent studies beginning to reveal their significant impact on algorithmic performance.
A comprehensive understanding of the origins of demographic accuracy differentials requires disentangling the complex interplay between demographic attributes and confounding behavioral or environmental variables. This necessitates the development of training and evaluation datasets that are not only demographically diverse but also annotated for a broader set of non-demographic characteristics potentially correlated with recognition performance.
To systematically analyze these influences, it is essential to design datasets that enable controlled isolation– analogous to evaluating a partial derivative– of individual or combined factors. This would allow researchers to measure the effect of specific attributes or conditions on the performance of a FR model or pipeline.
Advancing our understanding of these causal relationships is critical not only for identifying the root causes of observed biases but also for facilitating systematic and reproducible evaluation. Ultimately, understanding these underlying causes is a prerequisite for developing effective mitigation strategies and building FR systems that are both fair and accurate.
@@
Recognizing the increasing reliance on remote verification, the MdTF and DHS S&T introduced the Remote Identity Validation Technology Demonstration (RIVTD) initiative [^4] Remote Identity Validation Technology Demonstration (RIVTD). In addition to security, accuracy, and liveness detection requirements, this program also places particular emphasis on ensuring demographic fairness in such technologies. A recent study by Fatima et al. [139] investigated demographic fairness in RIdV technologies using statistical methods to analyze performance disparities. Their analysis of five commercial RIdV systems revealed that only two achieved equitable outcomes across demographic groups. Notably, higher FNMRs were observed among African cohorts and individuals with darker skin-tones. Such findings highlight the necessity of evaluating RIdV technologies across demographic groups to ensure equitable and fair performance.
@@
Complex Bias Factors (Intersectionality): The majority of research on mitigation of accuracy differentials in FR, as discussed in Sec V has focused on single demographic attributes, such as race, age, or gender. However, several studies have identified that combination or intersection of various demographic factors causes (or amplifies) bias in FR models (cf. Sec II), while few works, such as [62, 43] have addressed the issue of disparities caused due to multiple or combined attributes. Existing bias mitigation techniques typically target one demographic attribute at a time, achieving measurable improvements in fairness for that specific attribute. However, in most cases, it remains unclear whether such processing inadvertently introduces imbalances in other demographic attributes. For instance, enhancing fairness for gender-related bias may increase disparities linked to ethnicity and vice-versa. This highlights the need for systematic evaluations of the intersectionality of demographic factors, such as race and gender combined. Consequently, developing mitigation methods capable of addressing multiple demographic attributes simultaneously remains an open challenge.
@@
Noisy Labels: The assignment of demographic attributes such as race, ethnicity, and skin-tone in FR datasets typically involves discrete labeling into finite categories. Some attributes, such as race, are often self-reported. In many cases, race and ethnicity annotations may be derived from automatic classifiers or manual efforts. Automatic classifiers, predominantly based on deep learning, are likely to be susceptible to bias too; while manual annotations are prone to human judgment errors. Similarly, skin-tone is frequently categorized using scales like Fitzpatrick’s, which discretizes it into specific values, ignoring its continuous spectrum. This reliance on discrete labels introduces noise into the training data, as samples near category boundaries are often inaccurately labeled. Recent work [108], addressed this issue by using probabilistic weights (soft labels) for demographic information instead of utilizing rigid (categorical) labels. However, most existing methods overlook the issue of errors in training data.